3. Calculate the heat released and Ecombustion when 0.263 g of caffeine (C8H10O2N4) is burned in a bomb calorimeter. The calorimeter constant is 3.78 kJ/C. and the temperature rises 1.525 degrees Celsius.
Answer #1

1)
Q cal = Ccal*delta T
Q cal = 3.78*(1.525-0.0)
Q cal = 5.764 KJ
Answer: 5.76 KJ

2)
This heat is supplied by C8H10O2N4

Molar mass of C8H10O2N4,
MM = 8*MM(C) + 10*MM(H) + 2*MM(O) + 4*MM(N)
= 8*12.01 + 10*1.008 + 2*16.0 + 4*14.01
= 194.2 g/mol


mass(C8H10O2N4)= 0.263 g

use:
number of mol of C8H10O2N4,
n = mass of C8H10O2N4/molar mass of C8H10O2N4
=(0.263 g)/(1.942*10^2 g/mol)
= 1.354*10^-3 mol
delta H = -Qcal/number of mol
delta H = -5.764/1.354*10^-3
delta H = -4.257*10^3 KJ/mol
Answer: -4.26*10^3 KJ/mol